A system of two paint buckets connected by a light Two paint buckets connected
Answer #1

PE lost by M2 =12kg bucket will be M2gh. PE gained by M1 = 4kg will be M1gh. Net loss of PE will be:

(M2-M1)gh. Total KE as it hits the floor will be: (1/2)(M1 + M2)v^2

(Note they both will be moving at the same speed because the rope doesn't stretch.)

Thus conserv. of Energy gives us:

(M2-M1)gh = (1/2)(M1 + M2)v^2 or

v = sqrt[2(M2-M1)gh / (M1 + M2)].

Plug in M1 = 4, M2 = 12, g=9.8, h=2 and you will get

v = 4.43 m/s
Answer #2

SOLUTIION : 


Let v be speed when 12 kg bucket  hits the floor. 


Net P. E. Lost = K.E gained by both the masses.


(12 - 4) * g * 2 = 1/2 (12 + 4) v^2

=> 16 g = 8 v^2

=> v = sqrt(2g) 

=> v = sqrt(2 * 9.8)

=> v = speed at which bucket hits the floor =  4.427 m/s (ANSWER)

answered by: Tulsiram Garg